SOLVE THE EQUATION AND FIND THE VALUE OF X FOR THE FOLLOWING EQUATION (12x−1)(4x−1)(3x−1)(6x−1)=5

SOLVE THE  EQUATION AND FIND THE VALUE OF X FOR THE FOLLOWING EQUATION (12x−1)(4x−1)(3x−1)(6x−1)=5
 

Grade:12

1 Answers

Vikas TU
7 years ago
(12x−1)(4x−1)(3x−1)(6x−1)=5
mukltiplying each term with each other we get the final eqn. as:
  864x4-720x3+210x2-25x-4 = 0
From remainder theorem,
x = 0.5 satifies it.
Hence (x – ½ ) is the factor of the bi – quadratic polynomial.
On dividing by the divisor (x – 1/2) we get,
   432x3-144x2+33x+4 = 0 
Now from remainder theorem, 
x = -1/12 satifies it.
Therfore on dividing the cubic eqn. by (12x + 1) we get the quadratic as:
36x2-15x+4 = 0
Final result can be written as: => (12x+1) ( x- 1/2) (36x2-15x+4) 
as 36x2-15x+4 has the discriminantt is less than zero.